OpenStudy (anonymous):

3y=3/4x+5 how do i write in standard form?

OpenStudy (anonymous):

ax+by=c 3y=3/4x+5 subtract 3/4x on both sides 3/4x+3y=5 multiply 4 on both sides so you can eliminate the fraction therefore 3x+12y=20

OpenStudy (anonymous):

what part do i subtract 3/4 on both sides

OpenStudy (anonymous):

what i meant by that is you just move the 3/4x on the other side so it will look like ax+by=c

OpenStudy (anonymous):

can you draw that for me

OpenStudy (anonymous):

 it's not as accurate but hope you get it